An object vibrating with a frequency 850 Hz causes resonance in a tube of air when the shortest length of air column is 10.3 cm. Calculate the second resonant length, assuming the speed of sound in air is 340 m/s.
1
Expert's answer
2020-05-26T12:50:16-0400

The condition for a resonance in a tube of air is (see https://en.wikipedia.org/wiki/Acoustic_resonance#Closed_at_one_end):


"f = \dfrac{nv}{4L}"

where "n" is positive integer (resonance mode), "v = 340 m\/s" is the speed of sound in air and "L = 0.103 m" is the length of the tube.

For a given conditions the number of resonance mode is:


"n = 4fL/v = 4\cdot 850\cdot 0.103/340 = 1"

The second resonant length on the same frequency will be:


"L = \dfrac{2v}{4f} = \dfrac{v}{2f} = \dfrac{340}{2\cdot 0.103} \approx 1650 m"

Answer. L = 1650 m.